OpenStudy (anonymous):

What is the answer to my previous post just need the amawer so I can see where I havr gone wrong , I cnt reply because its not letting me so dont ask me questions on this post-x /3+5=7

OpenStudy (solomonzelman):

are you able to message me?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Waitni can reply

OpenStudy (solomonzelman):

but I see you are able to reply... :)

OpenStudy (solomonzelman):

so your question is: \(\large\color{black}{ \frac{-x}{3}+5=7 }\) ?

OpenStudy (anonymous):

Is letting me reply now , sorry for the bad english

OpenStudy (solomonzelman):

your English is fine... so lets proceed with a question

OpenStudy (anonymous):

Not mimus x just x

OpenStudy (anonymous):

Minus

OpenStudy (solomonzelman):

\(\large\color{black}{ \frac{x}{3}+5=7 }\) this?

OpenStudy (anonymous):

Ya

OpenStudy (solomonzelman):

subtract 5 from both sides. \(\large\color{black}{ \frac{x}{3}+5 \color{red}{-5}=7 \color{red}{-5} }\)

OpenStudy (solomonzelman):

what do you get after doing this, can you show me?

OpenStudy (anonymous):

2

OpenStudy (solomonzelman):

yes, you get: \(\large\color{black}{ \frac{x}{3}=2 }\)

OpenStudy (solomonzelman):

then multiply both sides times 3

OpenStudy (solomonzelman):

\(\large\color{black}{ \frac{x}{3} \color{blue}{\times 3}=2 \color{blue}{\times 3} }\)

OpenStudy (solomonzelman):

show me what do you get after performing this step?

OpenStudy (anonymous):

Just tell me the answer people have already showed me the method please thanks , I want to see where I have gone wrong thanks

OpenStudy (solomonzelman):

but if I wanted to check your work,. then I don't really see much work done by you. You haven't posted what you got, have you?

OpenStudy (solomonzelman):

also, I am not allowed to just handle you the answer, you know that....

OpenStudy (anonymous):

@SolomonZelman sorry for the late reply , it wasn't letting me reply yesterday all of a sudden again. I wanted to figure out where I wwwent wrong independantly without any help from any one else. Thanks for your time though , sir. I know that you are not allowed to just handle answers but as I had got help on how to do this question previously I didn't want to waste my time as I was in a hurry. Your method of how you explain things is very labour intensive so I suggest you try o explain things as simple as possible and as quick as possible. I am a fast learner so there is no difficulty in you explain things like this.

OpenStudy (anonymous):

went